OpenStudy (katielong):

Converting areas and volumes: How do you convert units for example 1metre^2 into millimetres^2 ? or 0.45mm^2 into m^2 Thanks

OpenStudy (anonymous):

1m^2 = 1,000,000 mm^2, you can start here.

OpenStudy (akashdeepdeb):

1) 1m2 = 1m * 1m [1m = 100cm] = 100cm * 100cm [1cm = 10mm or 1m = 1000mm] = 1000mm * 1000mm = 1000000mm2 2) 0.45mm2 = 0.45mm * 1mm [1mm = 0.1cm] = 0.045cm * 0.1cm [1cm = 0.01cm or 1mm = 0.001m] = 0.00045m * 0.001m = 0.00000045m2 Understood? :)

OpenStudy (katielong):

sort of, but how about converting 1cm^3 into m^3 is there a technique/method?

OpenStudy (anonymous):

 try to answer it